# Graded Projection Recursion (GPR): A Framework for Controlling Bit-Complexity of Algebraic Packing

### v3

@@ 9 Complexity and implementation model
In exact dyadic lattice semantics, Appendix  B.5 proves that the certified base-selection rule (Eq.  271 with the one-step upgrade Eq.  272) succeeds after at most one upgrade at every node, provided coefficient-lattice alignment is preserved (i.e., no mantissa bits are dropped below the local $\mathcal{L}_{\mathrm{coeff}}(\sigma)$ resolution). Therefore, within the declared exact-execution regime, the ClassicalMM branch in Algorithm  1 is unreachable and serves only as a correct-on-return guard for overflow/truncation (or implementation bugs). On fixed-width hardware, if a run ever leaves this regime, one may either take the conservative ClassicalMM fallback, or increase the staged-precision budget / word width and retry. Algorithm  2 formalizes a simple retry-with-more-staging wrapper, and Theorem  9.2 shows that it preserves the soft-quadratic bit/word complexity when the required precision $b(n)$ is polylogarithmic.
@@ 9.1 Retry with more staging (optional)
The ClassicalMM branch in Algorithm  1 is a correct-on-return backstop for implementations that leave the declared exact-execution regime (Appendix  B.5). In staged-precision fixed-word mode, a more efficient response is often to increase the staged mantissa budget and rerun the computation. The wrapper below does so without changing the kernel logic.
@@
Let KernelMM_Attempt(A,B;K)(A,B;K) denote Algorithm  1 executed in digit-band staged mode with $K$ words per scalar, except that the failure-policy fallback of Appendix  B.5 is replaced by returning a failure flag and aborting the attempt (i.e., no extracted value computed under a failed certificate is ever returned).
Input: $A,B\in\mathcal{L}_{D}^{n\times n}$ with $n$ a power of two; initial staging $K_{0}\geq 1$; maximum staging budget $K_{\max}$
Output: $C=AB$ exactly (or exact fallback if the staging limit is exhausted)
$K\leftarrow K_{0}$
@@
 $(C,\mathrm{ok})\leftarrow\textsc{KernelMM\_Attempt}(A,B;K)$
@@
    return $C$
 $K\leftarrow 2K$
return ClassicalMM(A,B)(A,B)
// correct-on-return backstop
[…]
@@
Fix a base word width $w$ and consider digit-band (slice) staging with $K$ words per scalar. Assume that for a full run of Algorithm  1 under $K$-word staging in the certified (no-failure) regime, the fixed-word cost satisfies $B(n,K)=\Theta(n^{2}\log n\cdot K)$ (each scalar primitive costs $\Theta(K)$ word operations).
@@
On the successful iteration, no node triggers the failure policy, so the attempted execution follows exactly the same computation path as Algorithm  1 and satisfies the hypotheses of Theorem  8.5, so it returns $C=AB$ exactly. If instead the wrapper exhausts the staging limit ($K>K_{\max}$), it returns ClassicalMM(A,B)(A,B), which is exact by Algorithm  3.
@@
$\sum_{i=0}^{s}B(n,K_{i})~=~\Theta(n^{2}\log n)\sum_{i=0}^{s}K_{i}~=~\Theta(n^{2}\log n)\,(2K_{s}-K_{0})~=~O(n^{2}\log n\cdot K_{s}).$ (74)
@@ 9.2 Arithmetic recurrence
    applying the sieve entrywise to extract the middle bands,
@@
    and performing the rank-one centering correction.
@@
Let $T(n)$ denote the arithmetic time of KernelMM on $n\times n$ inputs, measured as the number of scalar ring operations (additions, multiplications by fixed constants such as powers of two, and nearest-integer rounding calls), assuming $n$ is a power of two and that Appendix  B.5’s certificate-and-upgrade rule succeeds at every recursion node (so the algorithm never takes the ClassicalMM fallback), then:
@@
We now consider a single call of KernelMM on size $n$ under the assumption that the certificate-and-upgrade rule of Appendix  B.5 succeeds at this node and recursively at all descendants (so no fallback occurs).
@@ 9.3 Bit/word complexity (fixed words, staged precision)
Fix a word width $w$ for mantissa storage, and assume exponent offsets and loop indices fit in fixed-width machine integers. Assume additionally that Appendix  B.5’s certificate-and-upgrade rule succeeds at every recursion node (so the algorithm never takes the ClassicalMM fallback). Let $b(n)$ denote the required effective mantissa precision (in bits) needed to represent packed temporaries and preserve the extractor slack $\delta$ at all decision sites for dimension $n$ under the chosen execution discipline.
@@ 9.4 Example: integer polynomial convolution as a GPR-admissible packing
Therefore, choosing $\beta$ so that the master gap inequality holds coefficientwise (Lemma 5.3) the master gap inequality holds coefficientwise (Lemma 5.3) gives exact middle-band recovery coefficientwise by Lemma 5.2. For example, any $\beta$ satisfying:
@@ 10 Prior related work and what is new here
The core algebraic trick behind our quadrant packing is not new: embed multiple quantities into a single multiplication by introducing a large parameter (a base $\beta$, or a formal variable $\lambda$), and then decodes a desired coefficient/band by an extractor (rounding, digit extraction, interpolation, CRT, etc.). Variants of this idea appear across several communities.
@@ 13.1 Guarded dyadic $\ell_{2}$ normalization (float-GPR)
Lemma 13.9 (Sieve stability under additive drift).
Fix $\beta\geq 2$ and consider a packed scalar:
$z^{\star}~=~\beta U+T+L/\beta$ (123)
with $U,T\in D^{-2}\mathbb{Z}$. Let $\delta\in(0,1/2]$, and suppose the ideal master gap holds with slack $\delta$, i.e.:
$\lvert D^{2}L/\beta\rvert+\lvert D^{2}T/\beta\rvert+\lvert D^{2}L/\beta^{2}\rvert~\leq~\frac{1}{2}-\delta.$ (124)
Let the implemented scalar be $z=z^{\star}+\eta$ with additive drift $\eta\in\mathbb{R}$, then if:
$\lvert D^{2}\eta\rvert~\leq~\delta,$ (125)
the sieve extraction is invariant, i.e., $\mathcal{S}_{\beta}(z)=\mathcal{S}_{\beta}(z^{\star})=T$.
@@
The sieve uses only the two rounded quantities $\left\lfloor D^{2}z\right\rceil$ and $\left\lfloor D^{2}z/\beta\right\rceil$. Under the ideal slack $\delta$, both rounding arguments are at distance at least $\delta$ from their respective half-integer boundaries. The additive perturbation $\eta$ shifts each rounded argument by at most $\delta$ in the lifted units, so Lemma 12.10 (or Lemma 4.1 applied to the lifted arguments) implies the rounding outcomes do not change. Therefore the sieve output does not change. ∎
@@
    Recovering middle-band quantities via sieve-style extraction decisions (the same discrete decisions as in Part I), but with all primitive operations executed in format $\mathsf{F}$.
@@
    Ledger-bounded drift: The total drift between the implemented and ideal decision statistics is bounded by $\varepsilon_{\mathrm{tot}}\leq\delta$ (Definition 12.9).
@@ 17.2 APSP via min-plus products (semiring mode)
$X=\Delta_{X}\mathbf{1}\mathbf{1}^{\top}+R_{X},$ (179)
@@ 19 GPR-admissible packing: the master condition
If Equation 189 holds, then this is at most $1/2-\delta$, so Lemma 5.3 applies entrywise. Therefore the two sieve gap inequalities hold and Lemma 5.2 gives $\mathcal{S}_{\beta}(Z_{\ell})=T_{\ell}$ exactly for that entry. Because the argument is uniform in $\ell$ and in the entry, extraction is exact at every depth.
@@ 21.4 B) Beyond rounding gaps: an extractor abstraction
  • Rounding-gap extractor (Part I): Here $\mathsf{Packed}=\mathbb{R}$ (or a coefficient lattice), $\mathrm{Extract}=\mathcal{S}_{\beta}$, and $\mathsf{Cond}$ is the master gap inequality ensuring band separation.

### v4

@@ 9 Complexity and implementation model
In exact dyadic lattice semantics, Appendix B.5 proves that the certified base-selection rule (Eq. 296 with the one-step upgrade Eq. 297) succeeds after at most one upgrade at every node, provided coefficient-lattice alignment is preserved (i.e., no mantissa bits are dropped below the local $\mathcal{L}_{\mathrm{coeff}}(\sigma)$ resolution). Therefore, within the declared exact-execution regime, the ClassicalMM branch in Algorithm 1 is unreachable and serves only as a correct-on-return guard for overflow/truncation (or implementation bugs). On fixed-width hardware, if a run ever leaves this regime, one may either take the conservative ClassicalMM fallback, or increase the staged-precision budget / word width and retry. Algorithm 2 formalizes a simple retry-with-more-staging wrapper, and Theorem 9.2 shows that it preserves the soft-quadratic bit/word complexity when the required precision $b(n)$ is polylogarithmic.
@@ 9.1 Retry with more staging (optional)
The ClassicalMM branch in Algorithm 1 is a correct-on-return backstop for implementations that leave the declared exact-execution regime (Appendix B.5). In staged-precision fixed-word mode, a more efficient response is often to increase the staged mantissa budget and rerun the computation. The wrapper below does so without changing the kernel logic.
@@
Let KernelMM_Attempt(A,B;K)(A,B;K) denote Algorithm 1 executed in digit-band staged mode with $K$ words per scalar, except that the failure-policy fallback of Appendix B.5 is replaced by returning a failure flag and aborting the attempt (i.e., no extracted value computed under a failed certificate is ever returned).
[…]
$A,B\in\mathcal{L}_{D}^{n\times n}$ with $n$ a power of two; initial staging $K_{0}\geq 1$; maximum staging budget $K_{\max}$.
$C=AB$ exactly (or exact fallback if the staging limit is exhausted).
$K\leftarrow K_{0}$.
@@
  $(C,\mathrm{ok})\leftarrow\textsc{KernelMM\_Attempt}(A,B;K)$.
@@
   return $C$.
  end if
  $K\leftarrow 2K$.
end while
return ClassicalMM(A,B)(A,B) $\triangleright$ correct-on-return backstop.
@@
Fix a base word width $w$ and consider digit-band (slice) staging with $K$ words per scalar. Assume that for a full run of Algorithm 1 under $K$-word staging in the certified (no-failure) regime, the fixed-word cost satisfies $B(n,K)=\Theta(n^{2}\log n\cdot K)$ (each scalar primitive costs $\Theta(K)$ word operations).
@@
On the successful iteration, no node triggers the failure policy, so the attempted execution follows exactly the same computation path as Algorithm 1 and satisfies the hypotheses of Theorem 8.5, so it returns $C=AB$ exactly. If instead the wrapper exhausts the staging limit ($K>K_{\max}$), it returns ClassicalMM(A,B)(A,B), which is exact by Algorithm 4.
@@
$\sum_{i=0}^{s}B(n,K_{i})$ $=$ $\Theta(n^{2}\log n)\sum_{i=0}^{s}K_{i}$ (70)
$=$ $\Theta(n^{2}\log n)(2K_{s}-K_{0})$ (71)
$=$ $O(n^{2}\log n\cdot K_{s}).$ (72)
@@ 9.2 Arithmetic recurrence
    Applying the sieve entrywise to extract the middle bands, and
@@
    Performing the rank-one centering correction.
@@
Let $T(n)$ denote the arithmetic time of KernelMM on $n\times n$ inputs, measured as the number of scalar ring operations (additions, multiplications by fixed constants such as powers of two, and nearest-integer rounding calls), assuming $n$ is a power of two and that Appendix B.5’s certificate-and-upgrade rule succeeds at every recursion node (so the algorithm never takes the ClassicalMM fallback), then:
@@
We now consider a single call of KernelMM on size $n$ under the assumption that the certificate-and-upgrade rule of Appendix B.5 succeeds at this node and recursively at all descendants (so no fallback occurs).
@@ 9.3 Bit/word complexity (fixed words, staged precision)
Fix a word width $w$ for mantissa storage, and assume exponent offsets and loop indices fit in fixed-width machine integers. Assume additionally that Appendix B.5’s certificate-and-upgrade rule succeeds at every recursion node (so the algorithm never takes the ClassicalMM fallback). Let $b(n)$ denote the required effective mantissa precision (in bits) needed to represent packed temporaries and preserve the extractor slack $\delta$ at all decision sites for dimension $n$ under the chosen execution discipline.
@@ 9.4 Example: integer polynomial convolution as a GPR-admissible packing
Therefore, choosing $\beta$ so that the master gap inequality holds coefficientwise (Lemma 5.3) gives exact middle-band recovery coefficientwise by Lemma 5.2. For example, any $\beta$ satisfying:
@@ 10 Prior related work and what is new here
The core algebraic trick behind our quadrant packing is not new: embed multiple quantities into a single multiplication by introducing a large parameter (a base $\beta$, or a formal variable $\lambda$), and then decode a desired coefficient/band by an extractor (rounding, digit extraction, interpolation, CRT, etc.). Variants of this idea appear across several communities.
@@ 13.1 Guarded dyadic $\ell_{2}$ normalization (float-GPR)
Lemma 13.9 (Scaled sieve stability under additive drift).
Fix $\beta\geq 2$ and let $\sigma\geq 0$ be the node-local coefficient-lattice exponent (Definition 3.2). Consider a packed scalar:
$z^{\star}~=~\beta U+T+L/\beta,$ (121)
then with $L\in\mathbb{R}$:
$U,T\in\mathcal{L}_{\mathrm{coeff}}(\sigma)~=~D^{-2}2^{-\sigma}\mathbb{Z}.$ (122)
Let $\delta\in(0,1/2]$, and suppose the ideal scaled master gap holds with slack $\delta$, i.e.:
$\lvert D^{2}2^{\sigma}L/\beta\rvert+\lvert D^{2}2^{\sigma}T/\beta\rvert+\lvert D^{2}2^{\sigma}L/\beta^{2}\rvert~\leq~\frac{1}{2}-\delta.$ (123)
Let the implemented scalar be $z=z^{\star}+\eta$ with additive drift $\eta\in\mathbb{R}$. If:
$\lvert D^{2}2^{\sigma}\eta\rvert~\leq~\delta,$ (124)
then the scaled sieve extraction is invariant:
$\mathcal{S}_{\beta,\sigma}(z)=\mathcal{S}_{\beta,\sigma}(z^{\star})=T.$ (125)
@@
Let $M:=D^{2}2^{\sigma}$ be the lifted scaling at this decision site. The scaled sieve $\mathcal{S}_{\beta,\sigma}$ depends only on the two nearest-integer roundings $\left\lfloor Mz\right\rceil$ and $\left\lfloor Mz/\beta\right\rceil$. Under the ideal slack $\delta$, both lifted rounding arguments are at distance at least $\delta$ from their respective half-integer boundaries. The additive perturbation $\eta$ shifts each lifted argument by at most $\lvert M\eta\rvert\leq\delta$, so Lemma 12.10 (or Lemma 4.1 applied to the lifted arguments) implies the rounding outcomes do not change. Therefore the sieve output does not change. ∎
@@
    Recovering middle-band quantities via the scaled sieve extractor $\mathcal{S}_{\beta,\sigma}$ (Definition 5.1), where the node-local lattice exponent $\sigma$ is computed from the (float) normalized operands as in Definition 3.2; all primitive operations are executed in format $\mathsf{F}$.
@@
    Ledger-bounded drift: The total drift between the implemented and ideal decision statistics is bounded by $\varepsilon_{\mathrm{tot}}\leq\delta$ (Definition 12.9). For scaled-sieve sites, the relevant decision statistics are the lifted rounding arguments MzMz and $Mz/\beta$ with $M=D^{2}2^{\sigma}$ (Lemma 13.9).
@@ 17.2 APSP via min-plus products (semiring mode)
$X~=~\Delta_{X}\mathbf{1}\mathbf{1}^{\top}+R_{X},$ (179)
@@ 19 GPR-admissible packing: the master condition
If Equation 189 holds, then this is at most $1/2-\delta$, so Lemma 5.3 applies entrywise. Therefore the two sieve gap inequalities hold and Lemma 5.2 gives $\mathcal{S}_{\beta,\sigma_{\ell}}(Z_{\ell})=T_{\ell}$ exactly for that entry. Because the argument is uniform in $\ell$ and in the entry, extraction is exact at every depth.
@@ 21.4 B) Beyond rounding gaps: an extractor abstraction
  • Rounding-gap extractor (Part I): Here $\mathsf{Packed}=\mathbb{R}$ (or a coefficient lattice), $\mathrm{Extract}=\mathcal{S}_{\beta,\sigma}$ (with the node-local lattice exponent $\sigma$), and $\mathsf{Cond}$ is the scaled master gap inequality ensuring band separation.
@@
21.5 C) Worked example: a folded $(\beta,\gamma)$ packing for $2\times 2$ block matrix multiplication
This subsection gives a “template package” for a folded multigraded packing of the $2\times 2$ block recursion for matrix multiplication (or any bilinear kernel). The point of the fold is that a single packed kernel evaluation returns two target coefficients (via a constant-size multi-degree band), reducing the node fan-out from four packed products (Algorithm refalg:KernelMM, Step (3)) to two.
Definition 21.11 (Folded multigraded embedding, target band, and coefficient-projection extractor).
Let $M,N$ be abelian groups (or modules) and let $\mathcal{B}:M\times M\to N$ be a bilinear kernel (Definition 18.1). For blocked matrix multiplication on $(m\times m)$ blocks, one may take $M=N=\mathbb{R}^{m\times m}$ and $\mathcal{B}(X,Y)=XY$.
Fix a node of a $2\times 2$ block recursion, and write the node-local operands as:
$A$ $=$ $\begin{pmatrix}A_{11}&A_{12}\\ A_{21}&A_{22}\end{pmatrix}$, (252)
$B$ $=$ $\begin{pmatrix}B_{11}&B_{12}\\ B_{21}&B_{22}\end{pmatrix}$, (253)
where each block lies in $M$. The desired block outputs are:
$C_{ij}~=~\mathcal{B}(A_{i1},B_{1j})+\mathcal{B}(A_{i2},B_{2j})\in N,~~~i,j\in\{1,2\}.$ (254)
The setting is the bivariate graded extension $M[\beta^{\pm 1},\gamma]$ and $N[\beta^{\pm 1},\gamma]$ (a $t=2$ instance of Definition 21.5 with bases $(\beta_{1},\beta_{2})=(\beta,\gamma)$).
Folded embeddings.
Define two left embeddings (one per block-row), each using the $\beta$-window $\{0,-1\}$:
$X_{i}(\beta,\gamma)~:=~A_{i1}+\beta^{-1}A_{i2}\in M[\beta^{\pm 1},\gamma],~~~i\in\{1,2\}.$ (255)
Define a single folded right embedding that packs both block-columns into the $\gamma$-window $\{0,1\}$:
$Y(\beta,\gamma)~:=~(B_{11}+\beta B_{21})+\gamma(B_{12}+\beta B_{22})\in M[\beta^{\pm 1},\gamma].$ (256)
Define the two packed kernel outputs:
$Z_{i}(\beta,\gamma)~:=~\mathcal{B}\!\left(X_{i}(\beta,\gamma),Y(\beta,\gamma)\right)\in N[\beta^{\pm 1},\gamma],~~~i\in\{1,2\}.$ (257)
Target band and coefficient projection extractor.
Let the target multi-degree band be:
$E_{0}~:=~\{(0,0),(0,1)\}\subset\mathbb{Z}^{2},$ (258)
corresponding to degrees $\beta^{0}\gamma^{0}$ and $\beta^{0}\gamma^{1}$. For a packed value:
$Z(\beta,\gamma)~=~\sum_{p,q}\beta^{p}\gamma^{q}Z^{[(p,q)]},$ (259)
write $Z^{[(p,q)]}$ for the coefficient of $\beta^{p}\gamma^{q}$. Define the extractor (Definition 21.8) by exact coefficient projection:
$\mathrm{Extract}(Z)~:=~\left(Z^{[(0,0)]},~Z^{[(0,1)]}\right)\in N\times N.$ (260)
Lemma 21.12 (Folded coefficient identity (two outputs per packed product)).
For each $i\in\{1,2\}$, the target-band coefficients of $Z_{i}=\mathcal{B}(X_{i},Y)$ satisfy:
$Z_{i}^{[(0,0)]}$ $=$ $\mathcal{B}(A_{i1},B_{11})+\mathcal{B}(A_{i2},B_{21})~=~C_{i1},$ (261)
$Z_{i}^{[(0,1)]}$ $=$ $\mathcal{B}(A_{i1},B_{12})+\mathcal{B}(A_{i2},B_{22})~=~C_{i2},$ (262)
then $\mathrm{Extract}(Z_{i})=(C_{i1},C_{i2})$.
Proof.
Expand the embeddings by explicit $(\beta,\gamma)$ degrees:
$X_{i}$ $=$ $\beta^{0}\gamma^{0}A_{i1}+\beta^{-1}\gamma^{0}A_{i2},$ (263)
$Y$ $=$ $\beta^{0}\gamma^{0}B_{11}+\beta^{1}\gamma^{0}B_{21}+\beta^{0}\gamma^{1}B_{12}+\beta^{1}\gamma^{1}B_{22}.$ (264)
By bilinearity and the convolution identity for graded bilinear embeddings (Lemma 20.1) applied coefficientwise in the bivariate grading, the coefficient at degree $(p,q)$ is the sum of $\mathcal{B}(\cdot,\cdot)$ over pairs of input degrees adding to $(p,q)$.
For $(p,q)=(0,0)$, the contributing pairs are $(0,0)+(0,0)$ and (1,0)+(1,0)(-1,0)+(1,0), giving $\mathcal{B}(A_{i1},B_{11})+\mathcal{B}(A_{i2},B_{21})$.
For $(p,q)=(0,1)$, the contributing pairs are $(0,0)+(0,1)$ and (1,0)+(1,1)(-1,0)+(1,1), giving $\mathcal{B}(A_{i1},B_{12})+\mathcal{B}(A_{i2},B_{22})$. ∎
Algorithm 3 FoldedGPR(A,B)(A,B) (folded $(\beta,\gamma)$ node)
Node-local operands $A,B$ for a bilinear kernel $\mathcal{B}:M\times M\to N$; recursion depth $\ell$ implicit.
$C=\mathcal{B}(A,B)$ in $N$.
(1) Recursive invariant amplification.
[…]
if base case then
[…]
end if
(2) Folded multigraded embedding.
Partition $\widehat{A},\widehat{B}$ into $2\times 2$ blocks $(\widehat{A}_{ij})$, $(\widehat{B}_{ij})$.
for $i\leftarrow 1$ to 22 do
  $X_{i}\leftarrow\widehat{A}_{i1}+\beta^{-1}\widehat{A}_{i2}$.
end for
$Y\leftarrow(\widehat{B}_{11}+\beta\widehat{B}_{21})+\gamma(\widehat{B}_{12}+\beta\widehat{B}_{22})$.
(3) Two recursive kernel calls and multi-band extraction.
for $i\leftarrow 1$ to 22 do
  $Z_{i}\leftarrow\textsc{FoldedGPR}(X_{i},Y)$.
  $(C_{i1},C_{i2})\leftarrow\mathrm{Extract}(Z_{i})=(Z_{i}^{[(0,0)]},Z_{i}^{[(0,1)]})$.
end for
Assemble $\widehat{C}\leftarrow\begin{pmatrix}C_{11}&C_{12}\\ C_{21}&C_{22}\end{pmatrix}$.
Undo the projection on $\widehat{C}$ and return $C$.
Proposition 21.13 (BWBM/master-condition view, extractor correctness, and fan-out reduction).
Assume the node-local invariant amplification enforces the usual depth-$\ell$ classical bound $S_{\ell}$ for atomic kernel evaluations (Definition 18.2). Then the folded construction in Definition 21.11 is GPR-admissible for the target band $E_{0}=\{(0,0),(0,1)\}$ (Definition 21.6) with a depth-independent constant, and the coefficient-projection extractor is correct. Moreover, the folded node replaces the four packed products of the unfurled $2\times 2$ quadrant rule by two packed products per node.
Proof.
(A) BWBM certificate in the $\beta$-grade (slice-by-$\gamma$ view).
Write $Y(\beta,\gamma)=Y_{0}(\beta)+\gamma Y_{1}(\beta)$ where:
$Y_{0}(\beta)$ $=$ $B_{11}+\beta B_{21},$ (265)
$Y_{1}(\beta)$ $=$ $B_{12}+\beta B_{22}.$ (266)
For each fixed $\gamma$-slice $q\in\{0,1\}$, the pair:
$X_{i}(\beta)$ $=$ $A_{i1}+\beta^{-1}A_{i2},$ (267)
$Y_{q}(\beta)$ $=$ $B_{1,q+1}+\beta B_{2,q+1}$ (268)
has exponent windows $D=\{0,-1\}$ and $D^{\prime}=\{0,1\}$ (width $\Delta=1$), no mixing ($\Lambda=1$), and bounded convolution width: for the target $\beta$-degree $e_{0}=0$ there are exactly two contributing pairs $(0,0)$ and (1,1)(-1,1) so $k_{T}=2$, while for the lower degree $e_{0}-1=-1$ the only contributing pair is (1,0)(-1,0) so $k_{L}=1$. Thus BWBM (Definition 20.2) holds (per slice) with:
$(\Delta,\Lambda,k_{T},k_{L})~=~(1,1,2,1),$ (269)
and Lemma 20.3 yields the induced (univariate) master constants:
$C_{T}$ $=$ $k_{T}\Lambda^{2}=2,$ (270)
$C_{L}$ $=$ $k_{L}\Lambda^{2}=1.$ (271)
(B) Multigraded master condition for the folded target band $E_{0}$.
By Lemma 21.12, each target coefficient $Z_{i}^{[(0,0)]}$ and $Z_{i}^{[(0,1)]}$ is a sum of exactly two atomic bilinear evaluations at depth $\ell$. Under the depth-$\ell$ classical bound, this gives:
$\|Z_{i}^{[(0,0)]}\|_{N}$ $\leq$ $2S_{\ell},$ (272)
$\|Z_{i}^{[(0,1)]}\|_{N}$ $\leq$ $2S_{\ell}.$ (273)
Therefore, the entire target band $E_{0}$ obeys a depth-uniform multigraded master condition (Definition 21.6) with a constant factor $C=2$ under max-aggregation across degrees, and $C=4$ under sum-aggregation across degrees.
(C) Extractor correctness.
Here $\mathrm{Extract}$ is literal coefficient projection onto distinct degrees in the formal graded ring, so its sufficient condition (Definition 21.8) is purely representational (“degrees do not collide”) and holds identically in the symbolic model. Under any numeric realization (e.g., mixed-radix in one base), the depth-uniform master bound from (B) is exactly the hypothesis needed to choose global packing parameters so that the sufficient condition holds, and Theorem 21.10 then gives correctness at every depth.
(D) Why the fold gives two children.
In the unfurled quadrant rule used in Algorithm 1 (Step (3)), the packed pairs $(X_{ij},Y_{ij})$ are formed for all $(i,j)\in\{1,2\}^{2}$, requiring four recursive kernel calls to recover the four blocks $C_{ij}$.
In the folded scheme, a single packed product $Z_{i}=\mathcal{B}(X_{i},Y)$ returns two target coefficients, $\beta^{0}\gamma^{0}$ and $\beta^{0}\gamma^{1}$, which are exactly $(C_{i1},C_{i2})$ by Lemma 21.12. Therefore only two recursive kernel calls (for $i=1,2$) are needed per node. ∎
Remark 21.14 (Finite-window model (no formal symbols)).
The same construction can be expressed without explicit $\beta,\gamma$ symbols by working over the constant support window:
$W~=~\{(-1,0),(0,0),(1,0),(-1,1),(0,1),(1,1)\}$ (274)
and representing packed values as coefficient tables $\{Z^{[(p,q)]}\}_{(p,q)\in W}$. Multiplication is then truncated degree-convolution (Lemma 20.1), and the extractor is coordinate projection onto $E_{0}=\{(0,0),(0,1)\}$.
Remark 21.15 (Algebraic view (no explicit “bands”)).
Note that explicit “bands” can be abstracted away by working in a finite support graded representation and treating packed values as coefficient tables (degrees -11,0,++11) or the fixed multigraded window in the folded construction). From this perspective, the three-band identity is just degree-convolution, and “extraction” is the linear projection onto the target degree(s). The dyadic sieve implements this projection numerically on the dyadic lattice by enforcing a rounding gap. This clarifies that the kernel is algebraic (constant-width extension with a fixed projection), so classical MM-based algebraic reductions transfer at the arithmetic-operation level, while our earlier GPR reductions are needed to preserve bounded representations in the bit/word model.